# Spectraformer: A Unified Random Feature Framework for Transformer

### v3

@@ 3.2 Component functions
Table 2: Summary of research gap in kernelized attention. Rows are the component functions, columns are the weight matrices. For each cell, the top item is the cited work for this combination for typical kernel tasks, while the bottom item is the cited work for Transformer-based tasks. $\times$ indicates no prior work for the combination. Spectraformer explores all combinations listed here, excluding ‘Base’ due to exhaustive work as seen in the table.
TrigRF PosRF OPRF SADERF
Monte Carlo Base (Rahimi & Recht, 2007) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
ORF (Yu et al., 2016) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
SORF (Yu et al., 2016) $\times$ $\times$ $\times$
@@
Unit-cube QMC (Avron et al., 2016) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
MM (Shen et al., 2017) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
Quadrature SGQ (Dao et al., 2017) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
Learner FastFoodL (Yang et al., 2015) $\times$ $\times$ $\times$
(Chowdhury et al., 2022) (Chowdhury et al., 2022) $\times$ $\times$
3.3 Utility
@@
Spectraformer is a generic random feature framework that allows for the combination of any weight matrix with any component function provided that the weight matrix satisfies Definition 3.1 and the component function satisfies Definition 3.2. We note that Spectraformer is an experimental rather than a theoretical framework. Due to the large number of weight matrices and component functions, it is infeasible for the time being to derive a general mathematical proof for the bounds and convergence of Spectraformer. Therefore, we seek to validate its feasibility and show its ability to discover novel and performant combination via experimental results.
3.4 Adding New Code to Spectraformer
We now discuss how to add a new component function or weight matrix to the base code for further implementation. New component functions need to satisfy Definition 3.2 and new weight matrices need to satisfy Definition 3.1.
3.4.1 Component Function
  1. 1.
    Add the new component function f to src/models/component_functions.py, the arguments should include data (the input), and other optional parameters.
  2. 2.
    Import f and add a new entry to FastAttention.comp_functions[f_name] = f in src/models/attention_performer.py (line 176)
3.4.2 Weight Matrix
  1. 1.
    Add the new weight matrix w to src/models/weight_matrix_approx.py or src/models/weight_matrix_learner.py, the arguments should include nb_rows (number of rows), nb_cols (number of columns) and device.
  2. 2.
    Import w and add the if clause and $w$ function call in src/models/attention_performer.py (line 208)
4 Results
4.1 Experiment Details
Table 3: Experimental results of Spectraformer variants on the LRA benchmark, over five seeds. We report mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)). $L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three. Entries are ranked by $\mu$ accuracy. The previous RF SOTAs (SADERF-ORF (Likhosherstov et al., 2023) and OPRF-ORF (Likhosherstov et al., 2022)) are in bold. We note that via Spectraformer, we find various variants outperforming these previous RF SOTAs in terms of accuracy, with different combinations offering different trade-off between accuracy, time, and memory. Entries are sorted by mean accuracy.
Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
L T R $\mu$ L T R $\mu$ L T R $\mu$
OPRF-FastFoodL 37.55 (0.48) 64.41 (0.62) 77.70 (0.33) 59.89 1.07 2.07 2.12 1.75 0.86 1.72 1.68 1.42
OPRF-MM 38.08 (0.53) 60.40 (0.85) 81.09 (0.18) 59.86 0.68 1.26 1.24 1.06 1.36 2.71 2.56 2.21
PosRF-MM 37.06 (0.37) 61.87 (1.79) 80.58 (0.53) 59.84 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
OPRF-ORF 38.34 (0.22) 60.16 (0.79) 80.88 (0.17) 59.80 0.68 1.26 1.25 1.06 1.36 2.71 2.56 2.21
SADERF-QMC 37.37 (0.38) 61.14 (1.48) 80.84 (0.11) 59.78 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
PosRF-QMC 37.11 (0.09) 61.69 (0.96) 80.55 (0.13) 59.78 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
SADERF-MM 37.10 (0.22) 60.68 (1.88) 81.13 (0.17) 59.64 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
SADERF-ORF 37.10 (0.19) 60.39 (2.08) 81.05 (0.22) 59.51 0.68 1.25 1.28 1.07 1.44 2.86 2.69 2.33
OPRF-QMC 37.69 (0.62) 59.94 (0.59) 80.38 (0.49) 59.34 0.68 1.26 1.26 1.07 1.36 2.71 2.56 2.21
SADERF-SGQ 37.11 (0.21) 62.46 (0.54) 78.38 (0.25) 59.32 0.68 1.25 1.27 1.07 1.44 2.86 2.69 2.33
SADERF-FastFoodL 36.02 (1.38) 64.63 (0.18) 76.99 (0.61) 59.21 1.07 2.08 2.16 1.77 0.92 1.84 1.80 1.52
OPRF-SGQ 37.10 (0.23) 61.25 (0.54) 78.69 (0.54) 59.01 0.67 1.26 1.25 1.06 1.36 2.71 2.56 2.21
PosRF-ORF 34.35 (5.96) 60.30 (0.97) 80.45 (0.22) 58.37 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
PosRF-FastFoodL 33.46 (3.70) 64.65 (0.36) 76.95 (0.48) 58.35 1.02 1.98 2.03 1.68 0.79 1.57 1.53 1.30
SADERF-SORF 33.30 (0.98) 64.70 (0.36) 74.71 (1.90) 57.57 0.68 1.24 1.28 1.07 1.44 2.86 2.69 2.33
PosRF-SGQ 28.64 (7.54) 62.38 (0.53) 78.28 (0.20) 56.43 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
OPRF-SORF 27.91 (3.26) 64.76 (0.66) 75.92 (1.74) 56.20 0.67 1.26 1.24 1.06 1.36 2.71 2.56 2.21
PosRF-SORF 21.27 (6.65) 62.99 (0.40) 67.10 (1.11) 50.45 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
@@ 4.1 Experiment Details
The LRA covers tasks of different sequence lengths, difficulty, and objective and it is designed to evaluate efficient Transformers, of which Spectraformer belongs to. Specifically, we evaluate the models on three textual LRA tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013). We discuss the task characteristics in greater details in Section 4.2. All our models are experimented on NVIDIA V100 and 12 CPU with 20GB of memory. Our codebase is based on and includes the adapted or original implementation from Chen et al. (2021); Chowdhury et al. (2022); Liu et al. (2022); Choromanski et al. (2021); Likhosherstov et al. (2022; 2023); Wang et al. (2020); Xiong et al. (2021), Kitaev et al. (2020); Thomas et al. (2018); Zhou et al. (2021) (Apache License, Version 2.0).
@@
We run our experiments on 5 different seeds due to computational limitation. The parameters are identical to Chen et al. (2021) and chosen to limit the parameter count to account for the marked training time of multiple models. All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being their random feature component. Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from Chen et al. (2021). Our hyperparameters are available in Table 8. The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
Our result in Table 3 shows that the top mean accuracy models, ranked in descending order, are: OPRF-FastFoodL, OPRF-MM, PosRF-MM, OPRF-ORF, SADERF-QMC. Our mean accuracy result indicates that novel Spectraformer combinations have outperformed the highlighted previous SOTA, i.e., OPRF-ORF and SADERF-ORF. However, we are aware that mean accuracy is insufficient in offering a comprehensive picture. Therefore, we first proceed with analyzing the statistics of the tasks before analyzing the significance of these results.
4.2 Task Characteristics
@@
We first analyze the characteristics of the three textual classification LRA (Tay et al., 2021) tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013). We take into consideration three main characteristics:
@@
  • Number of output classes. ListOps is a task which inputs a sequence containing brackets and simple mathematical operators, with the output being a integer between 1 and 10, hence it has 10 classes. Text is a sentiment classification task on an IMDb review dataset with the class being either negative or positive, hence it has 2 classes. Retrieval inputs two sequences and the output is either False (meaning the sequences are not of the same document) or True (meaning the sequences are of the same document).
@@
  • Maximum sequence size. ListOps is 2K; Text is 4K; whilst Retrieval requires the processing of two sequences each of 4K length making the total maximum sequence size being 8K.
@@
  • Required attention span. This is the ‘mean distance between the query token and the attended tokens, scaled by attention weights’. The average values for required attention span calculated by Tay et al. (2021) is: ListOps being 0.8, Text being 0.3, and Retrieval being 1.3.
@@
Table 4: Mean Pearson correlation between model performance statistics and each task characteristic across weight matrices and component functions. Model performance statistics include ‘acc’ (accuracy), ‘time’ (training time), ‘mem’ (peak memory consumption). Task characteristics are ‘size’ (maximum sequence size), ‘span’ (required attention span), and ‘classes’ (number of output classes). Correlation is calculated for each weight matrix or component function and the average value over weight matrix or component function respectively displayed. Left is weight matrix $W$, right is component function $f$.
$W$ acc time mem
size 0.93 0.74 0.64
span 0.37 0.02 -0.10
classes -0.91 -0.96 -0.94
$f$ acc time mem
size 0.93 0.56 0.60
span 0.36 0.02 -0.10
classes -0.91 -0.72 -0.89
We then calculate Pearson correlation of accuracy, training time, and memory consumption against these characteristics. Our result is shown in Table 4. The table indicates that: (1) contrary to intuition, longer sequence size gives higher accuracy but this is not because of increasing required attention span and they somewhat increase training time and peak memory consumption; (2) increasing the number of classes decreases accuracy, training time, and peak memory consumption.
4.3 Analysis of Combinations
Table 5: Mean statistics across weight matrices for each task. Task is denoted on the top left cell of each table. Mean statistics include ‘acc’ (accuracy (%)), ‘time’ (training time (hour)), and ‘mem’ (peak memory consumption (GB)).
ListOps acc time mem
SORF 27.49 0.64 1.32
SGQ 34.28 0.64 1.32
FastFoodL 35.68 1.05 0.86
ORF 36.59 0.64 1.32
QMC 37.39 0.64 1.32
MM 37.41 0.64 1.32
Text acc time mem
ORF 60.28 1.19 2.63
QMC 60.92 1.19 2.63
MM 60.98 1.19 2.63
SGQ 62.03 1.19 2.63
SORF 64.15 1.18 2.63
FastFoodL 64.56 2.04 1.71
Retrieval acc time mem
SORF 72.58 1.19 2.45
FastFoodL 77.22 2.10 1.67
SGQ 78.45 1.19 2.45
QMC 80.59 1.20 2.45
ORF 80.79 1.20 2.45
MM 80.93 1.20 2.45
We now explore the suitability of weight matrices and component functions depending on task characteristics. Table 5 shows the mean statistics across weight matrices for each task. We notice that weight matrices form two different groups which align with the required attention span differences. ListOps and Retrieval are tasks with high required attention span, ListOps being 0.8 and Retrieval being 1.3. In both these tasks, we observe that ORF, QMC, and MM perform comparatively against each other and a lot better against SORF, SGQ, and FastFoodL. However, the Text task, which has a much lower required attention span of 0.3, has much higher performance from the SORF-SGQ-FastFoodL group, whilst ORF, QMC, and MM perform much poorly on this task. Based on this observation, we can conclude that ORF, QMC, and MM are suited for high required attention span; whilst SORF, SGQ, and FastFoodL are suited for low required attention span. Analyzing mean statistics per task across component functions do not show any task-specific trend (refer to Table 9 for more information).
Table 6: Summary statistics of component functions $f$ (top) and weight matrices $W$ (bottom) in mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)). $f$ and $W$ entries are sorted by mean accuracy respectively.
Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
$f$ SADERF 59.17 1.19 2.19
$f$ OPRF 59.02 1.18 2.08
$f$ PosRF 57.20 1.02 1.77
$W$ MM 59.78 1.01 2.13
$W$ QMC 59.63 1.01 2.13
$W$ ORF 59.22 1.01 2.13
$W$ FastFoodL 59.15 1.73 1.41
$W$ SGQ 58.25 1.01 2.13
$W$ SORF 54.74 1.00 2.13

### v5

@@ 3.2.7. SADERF
Table 3. A summary of research gap in the literature on kernelized attention. The table shows combinations of weight matrices (columns) and component functions (rows). For each cell, the top item is the cited work for this combination for typical kernel tasks, while the bottom item is the cited work for Transformer-based tasks. $\times$ indicates that a combination has not been previously studied in either general kernel tasks or specifically for Transformers. Spectraformer explores all listed combinations (excluding ‘Base’ due to exhasutive pre-existing work)
$f$ Base ORF SORF QMC $M$ SGQ FastFoodL
TrigRF
(Rahimi and Recht, 2007)
(Choromanski et al., 2021)
(Yu et al., 2016)
(Choromanski et al., 2021)
(Yu et al., 2016)
$\times$
(Avron et al., 2016)
$\times$
(Shen et al., 2017)
$\times$
(Dao et al., 2017)
$\times$
(Yang et al., 2015)
(Chowdhury et al., 2022)
PosRF
(Choromanski et al., 2021)
(Choromanski et al., 2021)
(Choromanski et al., 2021)
(Choromanski et al., 2021)
@@
$\times$
(Chowdhury et al., 2022)
OPRF
(Likhosherstov et al., 2022)
(Likhosherstov et al., 2022)
(Likhosherstov et al., 2022)
(Likhosherstov et al., 2022)
$\times$ $\times$ $\times$ $\times$ $\times$
SADERF
(Likhosherstov et al., 2023)
(Likhosherstov et al., 2023)
(Likhosherstov et al., 2023)
(Likhosherstov et al., 2023)
$\times$ $\times$ $\times$ $\times$ $\times$
3.3. Utility
@@
Spectraformer is a generic random feature framework that allows for the combination of any weight matrix with any component function provided that the weight matrix satisfies Definition 3.1 and the component function satisfies Definition 3.2. Instructions on adding new component functions or weight matrices to Spectraformer are detailed step by step in the repository linked in the abstract. We note that Spectraformer is an experimental rather than a theoretical framework. Aside from the challenge of generalizing the large number of weight matrices and component functions, estimating the “actual error of the approximation, or how this error will propagate into downstream learning tasks” is not possible (Yao et al., 2023). This is because, as Yao et al. (2023) pointed out: the RFF literature has largely use theoretical bounds which are largely impractical due to being either highly conservative or making use of “unknown qualities” as evident in Section 3. Thus, we seek to validate its feasibility and show its ability to discover novel and performant combination via empirical results and statistical analyses.
4. Results
4.1. Experimental Details
@@
The LRA covers tasks of different sequence lengths, difficulty, and objective, and is designed to evaluate efficient Transformers. The LRA is preferred over other benchmarks such as GLUE (Linzen et al., 2018) since efficient Transformers are designed to address tasks with long input length which reduces Transformer’s training and inference speed to the default attention’s quadratic time complexity. The LRA consists of ListOps (Nangia and Bowman, 2018) (testing the parsing ability of models), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard and Ruder, 2018), Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013), Image (Image Classification on Sequence of Pixels in 1D projected from the 2D of images in the CIFAR-10 dataset (Krizhevsky et al., 2009)), and Pathfinder (Long-Range Spatial Dependency) (Linsley et al., 2018; Kim et al., [n. d.]). We discuss the task characteristics in greater details in Section 4.3. All our models are experimented on NVIDIA V100 and 12 CPU with 20GB of memory. Our codebase is based on and includes the adapted or original implementation from (Chen et al., 2021; Chowdhury et al., 2022; Liu et al., 2022; Choromanski et al., 2021; Likhosherstov et al., 2022, 2023; Wang et al., 2020; Xiong et al., 2021), (Kitaev et al., 2020; Thomas et al., 2018; Zhou et al., 2021) (Apache License, Version 2.0).
@@
The parameters are identical to (Chen et al., 2021) and chosen to limit the parameter count to account for the marked training time of multiple models. All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being the number of random feature. We vary the number of random feature between 64, 128, 256, and 512 for comprehensive analysis. Dimension as 128 was used in (Chen et al., 2021) and 256 was used in (Choromanski et al., 2021; Likhosherstov et al., 2022, 2023). Our hyperparameters are available in Table 6. Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from (Chen et al., 2021). The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
4.2. Overall Results
Our comprehensive experiments on the Long Range Arena (LRA) benchmark reveal that the Spectraformer framework not only produces highly competitive models but also establishes a new state-of-the-art for random feature-based efficient Transformers. Before diving into a detailed breakdown, we first present a high-level comparison of our top-performing variants against a wide range of established efficient Transformers in Table 1. This includes methods based on downsampling (Nystromformer (Xiong et al., 2021), Skyformer (Chen et al., 2021), BigBird (Zaheer et al., 2020)), learnable patterns (Reformer (Kitaev et al., 2020)), and other linearization approaches (Linformer (Wang et al., 2020)).
As shown in Table 1, our best model, OPRF-FastFoodL, achieves a mean accuracy of 57.17%, which is highly competitive with leading methods like Nystromformer (Xiong et al., 2021) (58.06%) and BigBird (Zaheer et al., 2020) (57.25%). This is a significant finding, as it demonstrates for the first time that a well-configured kernelized attention model can perform at the same level as other classes of efficient Transformers on this challenging benchmark. Notably, OPRF-FastFoodL is on par with BigBird (Zaheer et al., 2020) on mean accuracy while offering a substantial reduction in peak memory consumption (4.65 to 1.82 GB) and training time (3.32 hours to 2.26 hours)
To provide full transparency and detail on how our top variants were identified, Table 7 presents the complete results for all 18 combinations of component functions and weight matrices explored within Spectraformer, tested with random feature dimensions of 64, 128, 256, and 512. The previous state-of-the-art random feature models, OPRF-ORF and SADERF-ORF, are highlighted in bold for direct comparison.
The results in Table 8 confirm that several novel combinations discovered through our framework, such as OPRF-FastFoodL and PosRF-MM, consistently outperform the previous random feature SOTAs. To help navigate these extensive results, Table 4 summarises the best-performing model for each task according to accuracy, training time, and memory usage. This highlights the framework’s ability to generate variants that offer different trade-offs, catering to specific computational or performance requirements. We now proceed to a deeper analysis of these results.
Table 4. A summary of the best-performing Spectraformer models, with tasks as columns and optimization categories as rows. Results are shown for 64, 128, 256, and 512 feature settings.
Metric Feat Task
L T R I P $\mu$
Accuracy 64 OPRF-QMC OPRF-SORF OPRF-$M$ SADERF-FastFoodL PosRF-QMC OPRF-FastFoodL
128 OPRF-$M$ SADERF-SORF OPRF-ORF OPRF-FastFoodL OPRF-SGQ OPRF-FastFoodL
256 OPRF-FastFoodL SADERF-SORF SADERF-ORF OPRF-FastFoodL OPRF-FastFoodL OPRF-FastFoodL
512 OPRF-FastFoodL SADERF-SORF OPRF-$M$ OPRF-FastFoodL PosRF-ORF OPRF-FastFoodL
Time 64 PosRF-$M$ PosRF-QMC PosRF-SGQ PosRF-$M$ PosRF-$M$ PosRF-$M$
128 PosRF-SORF PosRF-QMC PosRF-QMC PosRF-QMC PosRF-SORF PosRF-QMC
256 PosRF-$M$ PosRF-SORF PosRF-$M$ PosRF-SGQ PosRF-QMC PosRF-$M$
512 PosRF-FastFoodL
Memory 64 PosRF-$M$
128 PosRF-FastFoodL
256 PosRF-FastFoodL
512 PosRF-FastFoodL
4.3. Task Characteristics
@@
The tasks in the LRA include: ListOps (Nangia and Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard and Ruder, 2018), Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013), Image (Image Classification), and Pathfinder (Long-Range Spatial Dependency) (Linsley et al., 2018; Kim et al., [n. d.]). We take into consideration three main characteristics:
@@
  • Number of output classes. ListOps is a task which inputs a sequence containing brackets and simple mathematical operators, with the output being a integer between 1 and 10, hence it has 10 classes. Text is a sentiment classification task on an IMDb review dataset with the class being either negative or positive, hence it has 2 classes. Retrieval inputs two sequences and the output is either False (meaning the sequences are not of the same document) or True (meaning the sequences are of the same document). Image is based on the CIFAR-10  (Krizhevsky et al., 2009) and has 10 classes. Pathfinder (Linsley et al., 2018; Kim et al., [n. d.]) has 2 classes and requires the model to identify the target contour between one end and the other.
@@
  • Maximum sequence size. ListOps is 2K; Text is 4K; whilst Retrieval requires the processing of two sequences each of 4K length making the total maximum sequence size being 8K; Image flattens the 2D of CIFAR-10 images of 32-by-32 into a list with the size of 1024; Similary, Pathfinder also takes in a 1024-dimensional list from 32-by-32 images.
@@
  • Required attention span. This is the ‘mean distance between the query token and the attended tokens, scaled by attention weights’. The average values for required attention span calculated by (Tay et al., 2021) is: ListOps being 0.7589, Text being 0.3194, Retrieval being 1.329, Image being 0.359 Pathfinder being 0.5371.
4.4. Analysis of Methods
Figure 4. Heatmaps of Spectraformer combinations for the 64, 128, 256, 512 features. $L$ and II have 10 classes, opposed to $T$, $R$, $P$ having 2 classes. Hence we adjust the vmin-vmax for the first group as 20-40 with the second group as 50-85.
@@ 4.4. Analysis of Methods
Table 5. Pearson correlation coefficients between model performance metrics (Mtr.) (Accuracy (Acc), Time (Time), Memory (Mem)) and task characteristics. Task characteristics include the number of classes (Cl.), input sequence length (Len.), and required attention span (Att.). The analysis is presented for component functions (left) and weight matrices (right).
$f$ $W$
Mtr. T PosRF OPRF SADERF FastFoodL QMC $M$ ORF SGQ SORF
Acc Cl. -0.908 -0.924 -0.921 -0.950 -0.921 -0.919 -0.921 -0.905 -0.913
Len. 0.706 0.720 0.722 0.664 0.735 0.741 0.728 0.754 0.696
Att. 0.459 0.516 0.494 0.404 0.566 0.562 0.557 0.506 0.356
Time Cl. 0.169 0.191 0.185 0.234 0.176 0.177 0.165 0.177 0.176
Len. -0.169 -0.198 -0.186 -0.215 -0.180 -0.180 -0.175 -0.184 -0.180
Att. -0.271 -0.275 -0.265 -0.364 -0.255 -0.255 -0.252 -0.259 -0.255
Mem Cl. 0.160 0.144 0.143 0.241 0.160 0.160 0.149 0.161 0.160
Len. -0.185 -0.157 -0.159 -0.253 -0.177 -0.177 -0.171 -0.181 -0.177
Att. -0.265 -0.233 -0.238 -0.399 -0.259 -0.259 -0.255 -0.263 -0.259
We now explore the suitability of weight matrices and component functions depending on task characteristics. Table 8 shows the mean statistics across component functions and weight matrices for each task as well as average performance across tasks. To ensure the performance of methods consistently across different feature settings, we also refer to Figure 4.
First, we will analyze the component functions and weight matrices in terms of accuracy. For component functions, we can see that OPRF and SADERF outperform PosRF for all tasks except Text. For weight matrices, we observe the following:
  • $T$: FastFoodL and SORF performs well in $T$ but other models also perform equally well in task $T$. This task is not too hard.
  • $R$: ORF, QMC, and MM perform the best. FastFoodL and SGQ perform well moderately, and SORF performs the worst.
  • $P$: Most models perform well (ORF, FastFoodL, QMC, MM). SGQ slightly less well than this group. SORF performs quite poorly.
  • $L$: QMC and MM perform the best. ORF performs moderately well. SORF performs the worst. Meanwhile, SGQ’s performance varies with respect to the number of random features.
  • II: FastFoodL performs the best, followed by SGQ. ORF, QMC and MM perform moderately. With SORF performing the worst.
Based on the above observations, we note that ORF, QMC, and MM tend to have the same trend. These weight matrices along FastFoodL (albeit having a different trend) have high mean performance (see Table 8) and high performance per task. This is followed by SGQ. SORF has the worst performance. QMC and MM having the same trend is expected since MM is an improvement on QMC using the same base formula and assumption. Empirically, QMC also outperforms ORF marginally as expected, based on existing work (see (Liu et al., 2022), Table 9 (Appendix) w.r.t. Gaussian training error). QMC and MM are based on Equation 3.1.5 which is a reformulation of Equation 9 whilst reducing variance. Thus, theoretically QMC and MM should behave similarly yet better than MC-based methods like ORF. SORF is the faster version of ORF (Yu et al., 2016), and empirically it becomes nearly unbiased with the number of features above 32, essentially collapsing to the bias of ORF. Yet in practice, Liu et al. (2022) show that SORF does not perform as well as ORF in classification tasks, and this is reflected in our experiments. In addition, Liu et al. (2022) also show that FastFoodF (which is the FastFood approximator (FastFood fixed) as opposed to FastFoodL, the FastFood learner) performs as well as QMC. Based on (Chowdhury et al., 2022) however, it is expected that FastFoodL exceeds this. And being the weight learner class, its behaviors deviate from the general trend of weight matrices. Liu et al. (2022) show empirically that GQ performs as well as QMC. However, GQ cannot be used in practice, hence the use of SGQ. It was also not compared empirically in (Liu et al., 2022). Given that SGQ approximates a kernel $k$ that is subgaussian with parameter $b$, $A\geq 24eb^{2}M^{2}$, $d$ being the dimension of the data, $M\subset\mathbf{R}^{d}$ be the diameter between inputs $x$ and $y$, then SGQ’s sampling complexity is bounded by $2^{d}\left(\frac{12eb^{2}M^{2}}{A}\right)^{A}$.
This means that the dimension of the token as well as the distance between two tokens being calculated will increase the complexity. Since our dimension of the token is quite high compared for most kernel tasks (64), and the distance between two tokens is quite high as well due to the tasks being from the LRA, this results in SGQ’s rather average performance. The relationship between the distance of the tokens and the approximation quality of SGQ is apparent in the moderate Pearson correlation coefficient of 0.506 between SGQ and attention span in Table 5. Unfortunately, we have less than an obvious idea for the relationship between the other weight matrices and token characteristics like SGQ since most of their sampling complexity is only dependent on the number of features (Liu et al., 2022).
Table 5 shows that, as expected, the number of classes is the primary determinant of models’ performance. However, the input length also plays a role in determining the outcome of the performance of a weight matrix in a task as well. We observe a similar trend across ORF, QMC, and MM as stated above across correlation coefficients between accuracy w.r.t. number of classes, input length, and attention span. SGQ as suggested above performs similarly for the first two metrics but with a slightly lower attention span coefficient. FastFood has the lowest correlation coefficients for both input length and attention span, suggesting that its learnable weights behave contrary to task characteristics and learning might be attributed to this.
Figure 5 shows the accuracy trend for different number of features for each task. We observe that accuracy peaks mostly at 128 and sometimes 256 number of features, thereafter maintaining or decreasingly stably. We also find that there is no relationship between input length and the number of features (also see Figure 6).
This suggests that increasing the number of random features past 128 induces overfitting and 128 is the ideal number of random features. The best way to confirm whether overfitting is occurring is to look at the gap between the generalization accuracy (the accuracy on the development set) and the training accuracy. We term this accuracy as the generalization gap. We calculate the average generalization gap and averaging them for all seeds of the same task and model between 512 - 256, 256 - 128, and 128 - 64 features. The result is respectively 0.143%, 0.041%, -0.404%. This confirms our initial intuition that 64 features are too few (underfitting) but 256 features are too many (overfitting). The accuracy obtained is obtained with early stopping, accuracy obtained without early stopping can induce a much larger generalization gap difference. This confirms that overfitting occurs past 128 features. In addition, during experimentation, we also found that the 512 features is highly unstable due to gradient explosion from large matrix calculation leading to NaN values. This occurred for three seed (twice for PosRF-SGQ and once for SAERF-ORF).
SORF in particular performs poorly at 64 number of features with accuracy improving drastically at 128 number of features. This validates previous findings that in transformer applications, 128 features is the optimal option.
Figure 5. Accuracy vs Number of Features Across All Tasks
4.5. Efficiency Analysis
Figure 6. Performance Metrics vs. Number of Features